A liquid with density of 59 lb / ft ^ 3 and viscosity of 7cp is transported at 105.4 gpm by a pum...

a liquid with density of 59 lb / ft ^ 3 and viscosity of 7cp is transported at 105.4 gpm by a pump 70% efficiency of a # 1 tank operating at 15psig to another tank # 2 that operates at 55psig. The level of liquid in tank # 2 is 86 ft above the liquid level in tank # 1. Between both tanks there is 350 ft of linear tube with internal diameter of 2.25in and e / D of 0.006; the following obstructions: two (2) angle valve open, two (2) globe valve open, three (3) return bends and five (5) square 90 ° elbows. Calculate: (a) the power of the pump (horsepower). (b) the daily operating cost if the energy costs $ 0.31 kW / hr.
Answer #1

Volumetric flow rale a = 10s. 4 gpm 1 gallon = 0-13368 ft3 3 -14.08 ft_ = 0.2348 -t- Sec 15 psig-21 604bt. Tressure in tank ǐk-> valus しリ k value fon two angle valve oPen ( value fon 2 globe wolue open i) k volue ton eun bunds K3- 15 K 1.3 -0.2348 Aofrom thu Moodys frichon faahon chent itorm friction fachon alcatn trichion losses ώ Head loss due to fuchen toht pipe! tube)4 Head loss olue to fricton for z nelunn bund 2 2X32.2 tS) Head loss loss due to friction fen S sjuane go* 2x32.2. hs = 7.37P=4497 kW b) claily opezatin^ cost Cost: 031 kw hi f0H 46-9チKW cost per day kw 14.65 x 24 Cost= 349.4 kw- Asue